If tickets for a show cost$2.00 for adults and $1.50 for children, how many of each kind of ticket were sold if a total of 300 tickets were sold for $525?
Answer by JulietG(1812)

If all 300 tickets were sold at adult prices, it would be $600.
The difference between the all-adult scenario and actual is 600-525, or $75
The difference in price between the adult and child ticket is .50.
$75 / .50 = 150
150 tickets were sold as child tickets.
That leaves 300-150=150 as adult tickets.
.
150 * 1.50 = $225
150 * 2.00 = $300
Total ticket sales - $525